OpenStudy (anonymous):

What is the solution to this inequality? x - 7 < 1

OpenStudy (welshfella):

add 7 to both sides of the inequality sign

OpenStudy (welshfella):

x - 7 + 7 = 1 + 7

OpenStudy (welshfella):

the -7 + 7 = 0 which leaves x on its own on the left side and 1 + 7 = ?

OpenStudy (welshfella):

* typo in my second post:- its < not =

OpenStudy (welshfella):

so we have x < 1 + 7